OpenStudy (anonymous):

Consider m(x)=Sin^2(x). Investigate the derivative and show that it can be written as 2sinxcosx.

OpenStudy (larseighner):

write out sin^2(x) as sin(x)sin(x) take the derivative using the product rule: (sin(x)sin(x))' = sin'(x)sin(x) + sin(x)sin'(x) if you got this far, you know sin'(x) = cos(x) it is just very simple algebra from there.

OpenStudy (anonymous):

try the dummy approach. lets sin^2(x), take sin(x) = w and then we've got sin^2(x) = w^2.... well the derivative of w^2 is 2*w or 2w whatever, okay we should used the chain I mean we've got 2w and 2w= 2(sin x) then multiply it with the derivative of sin(x), which is cos(x). so the problem is solved. derivative of sin^2(x) is 2sin(x)cos(x) or to be more elaborate is also sin^2(x) - cos^2(x).

OpenStudy (anonymous):

Hmmm let's see: Given f(x) = sin^2(x) and you wanna prove that it's equivalent to 2sinxcosx using the derivative: Well the derivative of sin^2(x) is Using the product rule: sinx'sinx + sinxsinx' cosxsinx + sinxcosx cosx(2sinx) or just simply 2sinxcosx That's one way to solve it :D OR you can think of sinx as u: Let u = sinx So rewritten as u^2 The derivative is this is simply : 2uu' Now using the substitution method: 2uu' = 2sinxsinx' Since sinx' = cosx The answer would be 2sinxcos :)